OpenStudy (anonymous):

40% of what number is 14

OpenStudy (danjs):

\[\frac{ 40 }{ 100 }x = 14\]

OpenStudy (anonymous):

no it doesnt

OpenStudy (danjs):

What percent * x/100 of - multiply what number - variable is - equals

OpenStudy (danjs):

1400 = 40x

OpenStudy (danjs):

x = 35

OpenStudy (danjs):

40% of 35 is 14

sammixboo (sammixboo):

Use this formula \(\rm \color{blueviolet}{\dfrac{part}{whole}=\dfrac{percent}{100}}\) Where 40 is the percent and 14 is the part Plug in the values \(\rm \color{blueviolet}{\dfrac{14}{whole}=\dfrac{40}{100}}\) Do you know how to solve this?

sammixboo (sammixboo):

We are trying to solve for the whole

OpenStudy (danjs):

forty percent [40/100] of (*) a number (x) is (=) 14